A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment.

n=12​

p=0.35​

x less than or equals 4
Answer #1

Solution

Given that ,

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X 4) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4)

= (12 C 0) * 0.350 * (0.65)12 + (12 C 1) * 0.351 * (0.65)11 + (12 C 2) * 0.352 * (0.65)10 +

(12 C 3) * 0.353 * (0.65)9 + (12 C 4) * 0.354 * (0.65)8

= 0.5833

Probability = 0.5833